OpenStudy (anonymous):

What is the solution for x in the proportion ten over sixteen equals five over x? 4 10 20 8

Directrix (directrix):

This is the equation you want to solve: 10/16 = 5/x

Directrix (directrix):

10 and 16 have a common factor of 2 so 10/16 can be written as 5/8. That gives: 5/8 = 5/x --> Solve for x @mathnoobie10

OpenStudy (anonymous):

I dont understand.. o _o

Directrix (directrix):

Take the equation and cross multiply. 5/8 = 5/x 5*x = 5*8 x = ?

OpenStudy (anonymous):

o _o Can u dumb it down a bit more .. Im new at this.

Directrix (directrix):

5 times what number is equal to 5 times 8 ?